OpenStudy (anonymous):

A box contains 4 yellow tiles, 6 green tiles, and 10 purple tiles. Without looking, you draw out a tile and then draw out a second tile without returning the first tile. Find P(purple, then purple)

OpenStudy (anonymous):

how many tiles total?

OpenStudy (anonymous):

20 tiles

OpenStudy (anonymous):

how many purple?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

ok, good, so the probability the first tile is purple is \(\frac{10}{20}=\frac{1}{2}\)

OpenStudy (anonymous):

now you pull out a purple tile, and don't replace it how many tiles total now?

OpenStudy (anonymous):

19?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and out of those 19, how many left are purple?

OpenStudy (anonymous):

9

OpenStudy (anonymous):

ok good so the probability that the second tile chosen is purple, if you know the first tile chosen was purple, is \(\frac{9}{19}\)

OpenStudy (anonymous):

your final job is to multiply the two probabilities together, i.e. compute \[\frac{1}{2}\times \frac{9}{19}\]

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

yw